# PET2Rep: Towards Vision-Language Model-Drived Automated Radiology Report Generation for Positron Emission Tomography

### v1

@@ Introduction
To assess the efficacy of VLMs in generating radiology reports, we evaluate generated results against the ground-truth reports with following aspects.
@@
In line with existing research, we adopt widely recognized natural language generation (NLG) evaluation metrics, including BLEU-n (Papineni et al. 2002), METEOR (Banerjee and Lavie 2005), and ROUGE-L (Lin 2004). Specifically, BLEU-n measures the similarity of word sequences by calculating the overlap of n-grams between the generated reports and reference reports. ROUGE-L conducts evaluation by comparing the longest common word subsequences, focusing on the alignment of texts. METEOR can effectively integrate synonyms and paraphrases, providing a more nuanced assessment of the semantic similarity between reports.
PET Clinical Efficiency Metrics.
NLG metrics primarily focus on word and sentence similarity while neglecting diagnostic accuracy. Reports with opposite diagnostic conclusions may achieve similar NLG scores. Conversely, results with correct uptake assessments but inconsistent formatting in the report text might receive lower NLG scores. Existing studies have explored the proposal of clinical efficacy (CE) metrics by utilizing text classifiers to extract abnormality labels for evaluation (Hamamci et al. 2024). However, these methods are not applicable to PET reports. To assess the clinical efficacy of PET reports, we introduce a series of clinical efficiency metrics to evaluate descriptions regarding radiotracer uptake patterns in key organs within generated PET reports. Given that the whole-body PET imaging data used in our study, we extract the assessment of uptake levels corresponding to each key organ from the report text and compare these assessments with the corresponding entries in the ground truth reports. For each key organ, we define four states of radiotracer uptake: Increased Uptake, Decreased Uptake, Absent Uptake, and Normal. Given the clinical focus on anomaly detection, we categorize the first three states into three distinct positive classes, with normal serving as the negative class. Our evaluation method involves independently calculating the precision, recall, and F1-score for each of the three positive classes. The final CE metrics are the macro-average across these three positive classes. The implementation details are elaborated in the Appendix. Compared to NLG metrics, CE metrics shift the evaluation from text-matching problem to multi-label classification assessment that more closely aligns with clinical diagnosis.
@@
Results and Analysis
Table. 1 summarizes the performance of the evaluated VLMs. After reviewing the evaluation results, we have drawn following conclusions.
Overall Ineffectiveness or Even Failure. All existing VLMs demonstrate limited effectiveness in the evaluation task. Shockingly, most VLMs even fail to surpass the performance of a simple template baseline. The requirement of a comprehensive, whole-body structured radiology report proved to be a significant hurdle for existing VLMs. However, we observe that most VLMs were incapable of consistently adhering to this structure. Some models produce chaotic, unusable, or even empty outputs, resulting in near-zero scores across all evaluation metrics, which are excluded from the table. Even when models attempted to follow the template, they frequently introduced irrelevant information or omitted mandatory sections, further highlighting their inability to grasp the task’s core requirements. This phenomenon suggests that many VLMs may be overfitted to the narrow data distributions of their training, which often focus on specific, limited tasks like single-organ interpretation or generic image captioning rather than holistic clinical reporting. Consequently, they fail to generalize to the complex requirements of clinical applications, where precision, completeness, and structural rigor are paramount.
State-of-the-Art Models Merely Match the Baseline. While the most advanced models like the Lingshu and MedGemma series, performed best relative to other VLMs, they only manage to achieve results barely comparable to the baseline. This underwhelming outcome indicates that even the leading edge of VLMs are still far from being ready for practical application in clinical workflow. While these models generate coherent text with high NLG metrics, their outputs often miss critical clinical details such as subtle tracer uptake abnormalities, resulting in low CE metrics. Further manual evaluation by radiologists demonstrate that the outputs of state-of-the-art models are also mostly unusable. The nuanced interpretation of tracer uptake patterns and the broad medical knowledge require for a whole-body assessment remain significant challenges, highlighting a critical gap between general language proficiency and specialized clinical expertise.
Increased Model Scale Does Not Guarantee Better Performance. Our evaluation also reveal an interesting phenomenon that within the same model series, larger-scale models do not necessarily perform better than smaller ones. In some cases, larger models appear to disregard the requirements of the task and generate some irrelevant content like fabricating patients’ names, ages and other information in the report, which ultimately hinders their evaluation performance. This counterintuitive finding suggests that for specialized and highly structured generation tasks like radiology report generation, architectural innovation and targeted training may be more critical than simply scaling up the model size.
More detailed analysis, clinical evaluation and case study of VLM-generated reports are elaborated in the Appendix.
Discussion and Conclusion
In this work, we present PET2Rep, the first comprehensive benchmark specifically designed for evaluating radiology report generation in PET imaging, addressing the critical gap between existing researches and clinical application. The benchmark consists of 565 carefully curated whole-body PET/CT image-report cases, representing a significant advancement in this domain. Another key innovation is the introduction of a series of clinical efficiency metrics to evaluate the quality of radiotracer uptake pattern description in key organs in generated reports PET reports, which is a decisive factor in clinical decision-making, as omissions in critical findings can alter therapeutic pathways.
Our experimental results starkly expose critical limitations in current VLMs. Despite their claimed strong performance on several multimodal medical benchmarks, these models universally fail to outperform even a simple template baseline in PET2Rep, with most producing chaotic or structurally non-compliant reports. These findings indicate that the field still demands fundamental innovations in clinically grounded evaluation frameworks and rigorous alignment with real-world reporting standards to attain true clinical viability. Many existing benchmark tasks assess the capabilities of VLMs through visual question answering for superficial understanding of the image, such as identifying the imaging modality or naming marked organs, which is far from the knowledge required for actual clinical diagnosis and treatment. While existing clinical-related report generation datasets mostly focus on local structural images, neglecting the integration of structural information and functional imaging for comprehensive whole-body evaluation. PET2Rep incorporates the rich experience of professional medical experts, providing crucial guidance for advancing large models toward general medical AI and guiding model development at a more refined level.
While PET2Rep represents a significant step forward, several limitations should be noted. Currently, our evaluations are restricted to 2D slices, which does not fully capture the 3D spatial relationships and volumetric information that are essential for comprehensive image interpretation. Furthermore, clinically important quantitative metrics such as standardized uptake value (SUV) and lesion volume measurements have been excluded from the current evaluation framework.
To address these limitations, we plan to expand the benchmark to incorporate full 3D PET/CT evaluations, enabling more comprehensive spatial and volumetric analysis. Critical quantitative measures including SUV and lesion volume will be reintegrated to enhance clinical validity. Additionally, while the benchmark currently supports only Chinese reports, we plan to extend it to multilingual evaluation, thereby improving generalizability and facilitating broader clinical adoption of VLMs across diverse healthcare settings. These planned enhancements will significantly improve the benchmark’s clinical relevance and utility for developing more robust PET report generation systems.
Appendix
A. Dataset Details
Data Collection.
Quality Control.
Data Visualization.
Figure 3: An example case with expert-annotated radiology report.
B. Evaluation Details
Natural Language Generation Metrics.
No. Organ (Anatomic Structures)
1 Cranium and Brain
2 Eyeballs
3 Nasal Cavity and Sinuses
4 Pharynx and Parapharyngeal Space
5 Palatine Tonsils and Larynx
6 Salivary Glands and Thyroid
7 Cervical Lymph Nodes
8 Lungs and Thoracic Cavity
9 Mediastinum and Heart
10 Esophagus
11 Liver
12 Gallbladder
13 Pancreas
14 Spleen
15 Kidneys and Adrenal Glands
16 Gastrointestinal Tract
17 Prostate/Uterus and Bladder
18 Abdominal and Pelvic Cavities
19 Spine and Bones
Table 2: List of key organs (anatomic structures) evaluated in PET Report clinical efficiency metrics.
PET Clinical Efficiency Metrics.
$\text{Precision}_{c}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}}$ (1)
$\text{Recall}_{c}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FN}_{c}}$ (2)
C. Preprocessing Details
CT Resampling and Normalization.
Figure 4: An example of CT image before and after preprocessing.
PET SUV Normalization.
Figure 5: An example of PET image before and after preprocessing.
Creation of Fused PET/CT Images.
Evaluation Dimension Score Meaning
4 Only minor inaccuracies, acceptable.
5 Completely accurate, highly consistent with image features.
2 Multiple important findings are omitted.
4 Findings are comprehensive, with only minor omissions.
2 Expression is unclear, with significant language issues.
2 Report has many issues, requires extensive revisions before use.
4 Basically usable, only requires minor polishing.
5 Can be directly used as a formal clinical report.
1 Overall quality is poor, unacceptable.
2 Has obvious problems, not recommended for use.
Overall Rating 3 Acceptable, but needs improvement.
4 Good performance, close to clinically usable standard.
5 High-quality generation, nearly flawless.
D. Additional Experiments and Discussion
Manual Evaluation by Human Experts.
To complement the automated quantitative metrics, we conduct a qualitative clinical evaluation to assess the practical utility and diagnostic reliability of the generated reports. For this assessment, we select the outputs from the highest-performing VLMs as determined by NLG and CE metrics. These selected model-generated reports alongside the corresponding ground truth reports are anonymized and presented to two radiologists for a blind review. The physicians are tasked with scoring each generated report based on the following dimensions, including Medical Accuracy (MedAcc), Key Findings Completeness (FinCom), Expression Clarity (ExpCla), Clinical Usability (CliUsa) and Overall Rating (OveRat). The scores range from 1 to 5 following the criteria in Table. 3, with higher scores indicating better performance. Manual evaluation results are shown in Figure. 6, where a larger and more outward-reaching polygon on the chart signifies higher scores across the evaluation dimensions. From the results, we can observe that moonshot-v1 is the best-performing model among the three VLMs for comparison. Comparing the two charts on the left (Expert A) with the two on the right (Expert B), the evaluation trends are highly consistent. However, none of the models have achieved scores that meet the stringent requirements of clinical practice, falling short of the thresholds necessary for reliable clinical application.
Figure 6: Performance comparison of three VLMs under different task settings for manual evaluation by two human experts rated across five dimensions.

### v2

@@ Introduction
To assess the performance of VLMs in radiology report generation, we compare the generated reports against the ground-truth reports using the following aspects.
Natural Language Generation (NLG) Metrics. In line with existing research, we adopt widely recognized NLG metrics, including BLEU-n (Papineni et al. 2002), METEOR (Banerjee and Lavie 2005), and ROUGE-L (Lin 2004). Specifically, BLEU‑n evaluates n‑gram overlap between generated and reference reports, ROUGE‑L measures alignment via the longest common subsequences, and METEOR accounts for synonyms and paraphrases to capture semantic similarity.
PET Clinical Efficacy (CE) Metrics. NLG metrics primarily focus on word and sentence similarity while neglecting diagnostic accuracy. Reports with opposite diagnostic conclusions may achieve similar NLG scores. Conversely, results with correct uptake assessments but inconsistent formatting in the report text might receive lower NLG scores. Existing studies have explored the proposal of clinical efficacy metrics by utilizing text classifiers to extract abnormality labels for CT report evaluation (Hamamci et al. 2024). However, these methods are not applicable to PET reports. To assess the clinical efficacy of PET reports, we introduce a series of CE metrics to evaluate descriptions regarding radiotracer uptake patterns in key organs within generated PET reports. Given that the whole-body PET imaging data used in our study, we extract the assessment of uptake levels corresponding to each key organ from the report text and compare these assessments with the corresponding entries in the ground truth reports. For each key organ, we define four states of radiotracer uptake: Increased Uptake, Decreased Uptake, Absent Uptake, and Normal. Given the clinical focus on anomaly detection, we categorize the first three states into three distinct positive classes, with Normal serving as the negative class. Our evaluation method involves independently calculating the precision, recall, and F1-score for each of the three positive classes. The final CE metrics are the macro-average across these three positive classes. The implementation details are elaborated in the Appendix. Compared to NLG metrics, CE metrics shift the evaluation from text-matching problem to multi-label classification assessment that more closely aligns with clinical diagnosis.
Model (year/month) NLG Metrics CE Metrics Overall
BL-4 MTR RG-L Pre Rec F1 (%)
Template Baseline 0.3150(0.0482) 0.1475(0.0141) 0.5110(0.0319) 0.2282(0.0179) 0.2220(0.0106) 0.2249(0.0123) 27.5
General-Purpose VLMs
Qwen2.5-VL-7B (25/1) 0.3050(0.0476) 0.1407(0.0198) 0.5075(0.0340) 0.2233(0.0236) 0.1974(0.0083) 0.2094(0.0132) 26.4
0.3057(0.0467) 0.1390(0.0186) 0.5088(0.0320) 0.2284(0.0227) 0.2023(0.0075) 0.2144(0.0121) 26.6
Qwen2.5-VL-32B (25/1) 0.1777(0.0421) 0.0063(0.0110) 0.4165(0.0516) 0.3402(0.0781) 0.0418(0.0127) 0.0743(0.0214) 17.6
0.1851(0.0408) 0.0063(0.0111) 0.4295(0.0486) 0.2728(0.0447) 0.0308(0.0047) 0.0554(0.0082) 16.3
Qwen2.5-VL-72B (25/1) 0.2223(0.0585) 0.0655(0.0172) 0.4234(0.0588) 0.2474(0.0513) 0.0295(0.0024) 0.0527(0.0043) 17.3
0.2273(0.0584) 0.0645(0.0171) 0.4306(0.0594) 0.2917(0.0328) 0.0393(0.0049) 0.0693(0.0084) 18.7
InternVL3-8B (25/4) 0.2439(0.0627) 0.0606(0.0443) 0.4739(0.0630) 0.2425(0.0151) 0.2107(0.0114) 0.2254(0.0119) 24.3
0.2509(0.0529) 0.0641(0.0463) 0.4845(0.0566) 0.2333(0.0153) 0.2099(0.0074) 0.2208(0.0087) 24.4
InternVL3-14B (25/4) 0.2513(0.0684) 0.0472(0.0528) 0.4835(0.0910) 0.2366(0.0206) 0.2057(0.0095) 0.2199(0.0129) 24.1
0.2495(0.0671) 0.0532(0.0506) 0.4813(0.0904) 0.2322(0.0196) 0.1982(0.0099) 0.2137(0.0131) 23.8
InternVL3-38B (25/4) 0.1377(0.0924) 0.0775(0.0483) 0.4371(0.1199) 0.2711(0.0203) 0.2072(0.0141) 0.2344(0.0127) 22.8
0.1446(0.0855) 0.0825(0.0480) 0.4618(0.0913) 0.2674(0.0258) 0.2435(0.0298) 0.2546(0.0278) 24.2
InternVL3-78B (25/4) 0.3090(0.0525) 0.1233(0.0359) 0.4997(0.0401) 0.2355(0.0255) 0.0520(0.0119) 0.0850(0.0157) 21.7
0.3090(0.0518) 0.1262(0.0318) 0.5008(0.0397) 0.2369(0.0492) 0.0748(0.0083) 0.1132(0.0122) 22.7
Yi-VL-6B (24/1) 0.0065(0.0316) 0.0002(0.0056) 0.0479(0.0709) 0.1144(0.0430) 0.0061(0.0020) 0.0115(0.0038) 3.1
0.0374(0.0733) 0.0029(0.0165) 0.1156(0.1432) 0.1519(0.0261) 0.0260(0.0033) 0.0444(0.0055) 6.3
Yi-VL-34B (24/1) 0.2610(0.1071) 0.0848(0.0664) 0.4439(0.1420) 0.2305(0.0159) 0.1869(0.0079) 0.2063(0.0098) 23.6
0.2854(0.0809) 0.0898(0.0645) 0.4779(0.0950) 0.2303(0.0211) 0.2038(0.0072) 0.2160(0.0116) 25.1
LLaVa-V1.5-7B (23/9) 0.1198(0.0508) 0.0126(0.0515) 0.3043(0.0639) 0.2044(0.0287) 0.1022(0.0091) 0.1306(0.0121) 14.6
0.0328(0.0141) 0.0056(0.0369) 0.1717(0.0283) 0.2460(0.0764) 0.0337(0.0093) 0.0592(0.0163) 9.2
OmniLMM-12B(24/4) 0.0412(0.0627) 0.0075(0.0232) 0.1339(0.1324) 0.1789(0.0330) 0.0173(0.0027) 0.0316(0.0050) 6.8
0.0397(0.0614) 0.0067(0.0238) 0.1293(0.1336) 0.2095(0.0393) 0.0180(0.0040) 0.0331(0.0071) 7.3
VisualGLM-6B 23/5 0.0361(0.0519) 0.0182(0.0517) 0.1338(0.1214) 0.0662(0.0710) 0.0002(0.0002) 0.0004(0.0004) 4.3
0.0306(0.0492) 0.0208(0.0588) 0.1173(0.1157) 0.3404(0.1494) 0.0014(0.0006) 0.0029(0.0012) 8.6
DeepSeek-VL2 (24/12) 0.2697(0.0675) 0.0939(0.0976) 0.4875(0.0536) 0.2170(0.0137) 0.1532(0.0076) 0.1795(0.0081) 23.4
0.2817(0.0637) 0.1054(0.0974) 0.4936(0.0476) 0.2198(0.0269) 0.1571(0.0135) 0.1831(0.0176) 24.0
Medical-Specific VLMs
MedDr(24/4) 0.2667(0.1012) 0.1564(0.0434) 0.4571(0.1168) 0.2270(0.0245) 0.1820(0.0201) 0.2020(0.0215) 24.9
0.2801(0.0874) 0.1536(0.0389) 0.4742(0.0951) 0.2397(0.0275) 0.2113(0.0084) 0.2243(0.0138) 26.4
HuatuoGPT-Vision (24/6) 0.1384(0.0865) 0.0000(0.0000) 0.3399(0.1112) 0.1692(0.0232) 0.0814(0.0186) 0.1097(0.0207) 14.0
0.2573(0.0546) 0.0743(0.0278) 0.4834(0.0577) 0.2183(0.0200) 0.1620(0.0148) 0.1859(0.0164) 23.0
MedVLM-R1 (25/2) 0.1602(0.1112) 0.0006(0.0097) 0.3472(0.1742) 0.2246(0.0285) 0.1019(0.0117) 0.1399(0.0150) 16.2
0.1708(0.1294) 0.0003(0.0070) 0.3358(0.1840) 0.2321(0.0324) 0.1204(0.0077) 0.1583(0.0110) 17.0
MedGemma-4B (25/7) 0.3015(0.0517) 0.1215(0.0384) 0.5077(0.0385) 0.2276(0.0185) 0.2260(0.0113) 0.2266(0.0129) 26.8
0.2874(0.0773) 0.1207(0.0339) 0.4875(0.0786) 0.2362(0.0162) 0.2245(0.0091) 0.2301(0.0103) 26.4
MedGemma-27B (25/7) 0.2185(0.0552) 0.0297(0.0157) 0.4390(0.0696) 0.2300(0.0375) 0.0391(0.0079) 0.0667(0.0130) 17.1
0.2251(0.0574) 0.0309(0.0153) 0.4521(0.0781) 0.2853(0.0435) 0.0846(0.0141) 0.1304(0.0201) 20.1
Lingshu-7B (25/6) 0.2848(0.0855) 0.1079(0.0727) 0.4793(0.0933) 0.2281(0.0162) 0.1970(0.0100) 0.2112(0.0097) 25.1
0.2775(0.0945) 0.1030(0.0748) 0.4700(0.1119) 0.2273(0.0220) 0.1942(0.0106) 0.2093(0.0138) 24.7
Lingshu-32B (25/6) 0.3050(0.0650) 0.1554(0.0520) 0.4999(0.0604) 0.2250(0.0178) 0.2035(0.0079) 0.2135(0.0109) 26.7
0.2987(0.0713) 0.1531(0.0597) 0.4939(0.0733) 0.2328(0.0151) 0.2071(0.0071) 0.2191(0.0091) 26.8
Closed-Source VLMs
Gemini 2.5 Pro (25/6) 0.1535(0.0411) 0.0186(0.0199) 0.3987(0.0438) 0.1705(0.0299) 0.0215(0.0056) 0.0381(0.0092) 13.4
0.1536(0.0420) 0.0199(0.0201) 0.4025(0.0477) 0.2394(0.0571) 0.0311(0.0066) 0.0550(0.0115) 15.0
GPT-4o (24/5) 0.2023(0.0422) 0.0287(0.0160) 0.4023(0.0421) 0.3375(0.0891) 0.0527(0.0110) 0.0910(0.0185) 18.6
0.2134(0.0425) 0.0318(0.0132) 0.4168(0.0412) 0.2540(0.0450) 0.0728(0.0085) 0.1130(0.0135) 18.5
Moonshot-v1 (25/1) 0.3064(0.0496) 0.1261(0.0301) 0.5157(0.0339) 0.2603(0.0220) 0.1457(0.0096) 0.1866(0.0117) 25.7
0.2923(0.0464) 0.1055(0.0302) 0.5142(0.0316) 0.2327(0.0232) 0.1803(0.0132) 0.2030(0.0160) 25.5
Qwen-VL-Max (25/1) 0.2315(0.0375) 0.0269(0.0035) 0.4462(0.0377) 0.2764(0.0313) 0.1897(0.0139) 0.2248(0.0183) 23.3
0.2479(0.0406) 0.0265(0.0047) 0.4649(0.0409) 0.2844(0.0220) 0.1802(0.0092) 0.2204(0.0113) 23.7
Figure 3: Performance comparison of three VLMs under different task settings for manual evaluation by two radiologists rated across five dimensions, including Medical Accuracy (MedAcc), Key Findings Completeness (FinCom), Expression Clarity (ExpCla), Clinical Usability (CliUsa) and Overall Rating (OveRat).
Results and Analysis
Table. 1 summarizes the performance of the evaluated VLMs, with the last column providing an overall reference score as the average of all metrics. After reviewing the evaluation results, we have drawn following conclusions.
Overall Ineffectiveness or Even Failure. All evaluated VLMs exhibit limited effectiveness in generating structured radiology reports. Alarmingly, most models fail to surpass even a simple template baseline. The requirement to produce comprehensive, whole-body structured reports presents a substantial challenge for existing VLMs. Many models are unable to consistently adhere to the prescribed report structure, occasionally generating disorganized, unusable, or even empty outputs, which yield near-zero scores across evaluation metrics and are thus omitted from the results table. Even when models attempt to follow the template, they often include irrelevant information or omit mandatory sections, underscoring their difficulty in capturing the core task requirements. This pattern suggests that many VLMs are overfitted to narrow training distributions, typically focused on specific tasks such as single-organ interpretation or generic image captioning, rather than holistic clinical reporting. As a result, they struggle to generalize to clinical applications, where accuracy, completeness, and structural consistency are essential.
State-of-the-Art Models Merely Match the Baseline. Although the most advanced models, such as the Lingshu and MedGemma series, outperform other VLMs, their performance remains only marginally comparable to the baseline. This underwhelming result indicates that even state-of-the-art VLMs are far from ready for practical application in clinical workflow. While these models can generate coherent text with high NLG metrics, they frequently omit critical clinical details, such as subtle tracer uptake abnormalities, leading to low CE metrics. Manual review by radiologists further confirms that the outputs of these models are largely unusable. The accurate interpretation of tracer uptake patterns, combined with the extensive medical knowledge required for comprehensive whole-body assessment, remains a major challenge, highlighting the gap between general language proficiency and specialized clinical expertise. Further manual evaluation by two radiologists in Figure. 3 demonstrate that the outputs of state-of-the-art models are also mostly unusable. The nuanced interpretation of tracer uptake patterns and the broad medical knowledge required for whole-body assessment remain significant challenges, highlighting a critical gap between general language proficiency and specialized clinical expertise.
Larger Model Does Not Necessarily Translate to Better Performance. Our evaluation reveals an intriguing phenomenon that within the same model series, larger-scale models do not consistently outperform their smaller counterparts. In some cases, larger models appear to overlook task requirements, generating irrelevant or fabricated details such as patient names and ages, which negatively impact their evaluation performance. This observation suggests that the inferior performance of larger models may not stem from model scaling itself, but rather from insufficient exposure to domain-specific data and task-oriented training. Therefore, for specialized and highly structured tasks like PET report generation, architectural innovation and targeted fine-tuning may play a more critical role.
Further details regarding experimental results analysis and case studies are presented in the Appendix.
Discussion and Conclusion
In this work, we present PET2Rep, the first comprehensive benchmark specifically designed for evaluating radiology report generation in PET imaging, addressing a critical gap between existing research and clinical application. The benchmark consists of 565 whole-body PET/CT image-report cases, representing a significant advancement in this domain. Another key innovation is the introduction of a series of clinical efficacy metrics to evaluate the quality of radiotracer uptake pattern description in key organs in generated reports PET reports, which is a decisive factor in clinical decision-making, as omissions in critical findings can alter therapeutic pathways.
Our experimental results clearly reveal the critical limitations of current VLMs. Despite their reported success on various multimodal medical benchmarks, all models fail to surpass even a simple template baseline in PET2Rep, with some models generating disorganized or structurally non-compliant reports. These findings underscore the need for fundamental advancements in clinically grounded evaluation frameworks and rigorous alignment with real-world reporting standards to achieve genuine clinical applicability. Many existing benchmark tasks assess the capabilities of VLMs through visual question answering, which primarily reflects superficial image understanding and falls short of the deep clinical reasoning required for diagnosis and treatment. Moreover, existing clinical report generation datasets are largely confined to localized anatomical structures, overlooking the integration of structural and functional information necessary for comprehensive whole-body evaluation. In this context, PET2Rep serves as an expert-informed and clinically aligned benchmark that helps bridge this gap, providing a foundation for exploring the potential of large models toward more generalizable medical intelligence and facilitating progress in domain-specific model development.
While PET2Rep represents a significant step forward, several limitations should be acknowledged. At present, our evaluations are limited to 2D slices, which do not fully capture the three-dimensional spatial relationships and volumetric information critical for comprehensive image interpretation (Zhang et al. 2022). Moreover, clinically important quantitative indicators, such as standardized uptake values (SUVs) and lesion volume measurements, are not yet incorporated into the current evaluation framework. To address these limitations, we plan to expand the benchmark to support full 3D PET/CT evaluations, enabling more complete spatial and volumetric analysis (Xue et al. 2025). Key quantitative measures, including SUVs and lesion volumes, will be reintegrated to enhance the benchmark’s clinical validity. In addition, while the current version supports only Chinese reports, future iterations will extend to multilingual evaluation, improving generalizability and facilitating broader clinical adoption of VLMs across diverse healthcare systems (Qiu et al. 2024). Another consideration lies in the contextual limitations of current VLMs. Generating long, structured whole-body reports may exceed the effective context length or reasoning capacity of some models. In future work, we plan to further investigate this issue by exploring hierarchical strategies, such as summarizing findings by anatomical regions followed by integration into a standardized report template. These planned enhancements will significantly improve the benchmark’s clinical relevance and utility for developing more robust report generation systems.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (Grant No. 82394432 and 92249302), and the Shanghai Municipal Science and Technology Major Project (Grant No. 2023SHZDZX02). The computations in this research were performed using the CFFF platform of Fudan University.
Appendix
A. Dataset Details
Data Collection.
Quality Control.
Data Visualization.
Figure 4: An example case with expert-annotated radiology report.
B. Evaluation Details
@@
No. Organ (Anatomic Structures)
1 Cranium and Brain
2 Eyeballs
3 Nasal Cavity and Sinuses
4 Pharynx and Parapharyngeal Space
5 Palatine Tonsils and Larynx
6 Salivary Glands and Thyroid
7 Cervical Lymph Nodes
8 Lungs and Thoracic Cavity
9 Mediastinum and Heart
10 Esophagus
11 Liver
12 Gallbladder
13 Pancreas
14 Spleen
15 Kidneys and Adrenal Glands
16 Gastrointestinal Tract
17 Prostate/Uterus and Bladder
18 Abdominal and Pelvic Cavities
19 Spine and Bones
Table 2: List of key organs (anatomic structures) evaluated in PET Report clinical efficacy metrics.
PET Clinical Efficacy Metrics.
$\text{Precision}_{c}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}}$ (1)
$\text{Recall}_{c}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FN}_{c}}$ (2)
C. Preprocessing Details
CT Resampling and Normalization.
Figure 5: An example of CT image before and after preprocessing.
PET SUV Normalization.
Figure 6: An example of PET image before and after preprocessing.
Creation of Fused PET/CT Images.
Evaluation Dimension Score Meaning
4 Only minor inaccuracies, acceptable.
5 Completely accurate, highly consistent with image features.
2 Multiple important findings are omitted.
4 Findings are comprehensive, with only minor omissions.
2 Expression is unclear, with significant language issues.
2 Report has many issues, requires extensive revisions before use.
4 Basically usable, only requires minor polishing.
5 Can be directly used as a formal clinical report.
1 Overall quality is poor, unacceptable.
2 Has obvious problems, not recommended for use.
Overall Rating 3 Acceptable, but needs improvement.
4 Good performance, close to clinically usable standard.
5 High-quality generation, nearly flawless.
D. Additional Experiments and Discussion
Manual Evaluation by Human Experts.
To complement the automated quantitative metrics, we conduct a qualitative clinical evaluation to assess the practical utility and diagnostic reliability of the generated reports. For this assessment, we select the outputs from the highest-performing VLMs as determined by NLG and CE metrics. These selected model-generated reports alongside the corresponding ground truth reports are anonymized and presented to two radiologists for a blind review. The physicians are tasked with scoring each generated report based on the following dimensions, including Medical Accuracy (MedAcc), Key Findings Completeness (FinCom), Expression Clarity (ExpCla), Clinical Usability (CliUsa) and Overall Rating (OveRat). The scores range from 1 to 5 following the criteria in Table. 3, with higher scores indicating better performance.
From manual evaluation results shown in Figure. 3, where a larger and more outward-reaching polygon on the chart signifies higher scores across the evaluation dimensions, we can observe that moonshot-v1 is the best-performing model among the three VLMs for comparison. Comparing the two charts on the left (Expert A) with the two on the right (Expert B), the evaluation trends are highly consistent. However, none of the models have achieved scores that meet the stringent requirements of clinical practice, falling short of the thresholds necessary for reliable clinical application.
@@ D. Additional Experiments and Discussion
Table 4: Detailed evaluation results of PET2Rep benchmark. Evaluation results presented in gray and white represent the results of separate PET and CT images and fused PET/CT images, respectively.